# On a discrete version of the position-momentum commutation relation

### v4

@@ VIII Discrete creation operator
where $\Psi_{0}(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\mathrm{Exp}[-\frac{1}{2}q^{2}]$ is the vacuum state. By following the analogy with the continuous case, we define
@@
Table 6: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain states $\psi_{C,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
IX Discrete version of the harmonic oscillator
The operator $\hat{\mathfrak{H}}\!$$\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\hat{\mathfrak{H}}\!=\!\frac{\hat{\mathfrak{p}}^{2}\!+\!\hat{\mathfrak{q}}^{2}}{2}$ (59)
@@ IX Discrete version of the harmonic oscillator
For any small $\varepsilon\!>\!0$, some of the eigenstates of $\hat{\mathfrak{H}}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 7), in the case d=31d\!=\!31, the eigenstates $\psi_{O,0}$, $\psi_{O,1}$, … , $\psi_{O,7}$ belong to $\mathcal{S}_{0.001}$. They form an orthonormal system.
Table 7: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied for certain states $\psi_{O,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
X Discrete coherent states
The discrete coherent states [2, 3] are
$\begin{array}[]{l}|n,k\rangle\!=\!\frac{1}{\sqrt{d}}D(n,k)|\mathbf{g}\rangle,\end{array}$ (61)
where $n,k\!\in\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$ and $D(n,k)\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\begin{array}[]{l}D(n,k)\psi(m)=\mathrm{Exp}[-\frac{\pi\mathrm{i}}{d}nk]\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}km]\,\psi(m\!-\!n),\end{array}$ (62)
@@ X Discrete coherent states
$\begin{array}[]{r}\parallel\!\mathfrak{C}|n,k\rangle\!\parallel^{2}=\!\!\!\sum\limits_{m=-s}^{s}\left|\frac{1}{d}\right.\!\!\sum\limits_{j=-s}^{s}\sum\limits_{\ell=-s}^{s}\!\!j(m\!-\!\ell)\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}j(m-\ell)]\\ \times\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k\ell]\,\mathbf{g}(\ell\!-\!n)\left.-\,\mathrm{i}\frac{d}{2\pi}\ \mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}km]\ \mathbf{g}(m\!-\!n)\right|^{2},\end{array}$ (63)
@@
$\parallel\!\mathfrak{C}|n,k\rangle\!\!\parallel=\parallel\!\mathfrak{C}|-n,k\rangle\!\!\parallel=\parallel\!\mathfrak{C}|n,-k\rangle\!\!\parallel=\parallel\!\mathfrak{C}|-n,-k\rangle\!\!\parallel.$ (64)
@@
Table 8: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain states $|j,k\rangle$.
$|n,k\rangle$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$ $|\ |\langle n,k|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$ $|n,k\rangle$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$ $|\ |\langle n,k|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$
@@
XI Discrete coherent state quantization
@@
$\begin{array}[]{l}g(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\,\mathrm{Exp}[-\frac{1}{2}q^{2}]\end{array}$ (66)
is the vacuum state, satisfy the resolution of the identity
@@
The map $f\!\mapsto A_{f}$ is called coherent state quantization [3, 15, 16] . One can prove that [15]
@@
by following the analogy with the continuous case, one can associate the operator $\mathfrak{A}_{f}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\begin{array}[]{l}\mathfrak{A}_{f}\!=\!\sum\limits_{n,k=-s}^{s}f(n,k)\,|n,k\rangle\langle n,k|\end{array}$ (71)
@@
defined on the discrete phase space. By using the discrete coherent state quantization $f\!\mapsto\!\mathfrak{A}_{f}$, we obtain for $f(n,k)\!=\!\frac{1}{2}(n^{2}\!+\!k^{2})$ the operator
$\begin{array}[]{l}\mathfrak{H}_{q}\!=\!-\frac{1}{2}\!+\!\sum\limits_{n,k=-s}^{s}\frac{n^{2}\!+\!k^{2}}{2}\,|n,k\rangle\langle n,k|\end{array}$ (73)
which can be regarded as a discrete version of the harmonic oscillator Hamiltonian. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{Q,n}$ of $\mathfrak{H}_{q}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 9), in the case d=31d\!=\!31, the states $\psi_{Q,0}$, $\psi_{Q,1}$, … , $\psi_{Q,7}$, where the states have been considered in the increasing order of eigenvalues, belong to $\mathcal{S}_{0.001}$.
Table 9: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain states $\psi_{Q,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
XII Discrete annihilation operator
It is known that, in the continuous-variable case, the standard coherent states $|q,p\rangle$ can be defined as eigenstates of the annihilation operator
By following the analogy, we consider the eigenstates $\psi_{A,n}$ of the discrete version $\hat{\mathfrak{a}}$ of $\hat{a}$. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{A,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 10), in the case d=31d\!=\!31, the eigenstates $\psi_{A,0}$, $\psi_{A,1}$, … , $\psi_{A,14}$, where the states have been considered in the increasing order of the number of sign oscillations, belong to $\mathcal{S}_{0.001}$.
h $\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$ $\psi_{A,0}$ $2.08125\!\times\!10^{-9}$ 0   $\psi_{A,1}$ $2.08125\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$ $\psi_{A,2}$ $2.08126\!\times\!10^{-9}$ $1.77636\!\times\!10^{-15}$ $\psi_{A,3}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $\psi_{A,4}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $\psi_{A,5}$ $1.41314\!\times\!10^{-7}$ $1.77636\!\times\!10^{-15}$ $\psi_{A,6}$ $1.41314\!\times\!10^{-7}$ 0   $\psi_{A,7}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$ $\psi_{A,8}$ $3.2588\!\times\!10^{-6}$ $2.66454\!\times\!10^{-14}$ $\psi_{A,9}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$ $\psi_{A,10}$ $3.2588\!\times\!10^{-6}$ $2.75335\!\times\!10^{-14}$ $\psi_{A,11}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,12}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,13}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,14}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,15}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,16}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,17}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,18}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,19}$ 0.0381896 $2.72907\!\times\!10^{-6}$
Table 10: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied by certain states $\psi_{A,k}$.
XIII Mehta states

### v5

@@ IV.2 Discrete version of Fock states
where $\Psi_{0}(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\mathrm{e}^{-\frac{1}{2}q^{2}}$ is the vacuum state. By following the analogy with the continuous case, we define
@@
Table 6: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation for certain states $\psi_{C,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
IV.3 Discrete version of the harmonic oscillator
The operator $\hat{\mathbf{H}}\!$$\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\hat{\mathbf{H}}\!=\!\frac{\hat{\mathbf{p}}^{2}\!+\!\hat{\mathbf{q}}^{2}}{2}$ (76)
@@ IV.3 Discrete version of the harmonic oscillator
For any small $\varepsilon\!>\!0$, some of the eigenstates of $\hat{\mathbf{H}}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 7), in the case d=31d\!=\!31, the eigenstates $\psi_{O,0}$, $\psi_{O,1}$, … , $\psi_{O,7}$ belong to $\mathcal{S}_{0.001}$. They form an orthonormal system. The relation $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx\mathrm{i}\frac{d}{2\pi}\psi$ is satisfied if and only if $\mathbf{a},\mathbf{a}^{\dagger}]\psi\approx\psi$. In this case $\frac{\hat{\mathbf{p}}^{2}\!+\!\hat{\mathbf{q}}^{2}}{2}\psi\!\approx\!(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\psi$, and the eigenvalues of $\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2}$ tend to be distributed equidistant,
$(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\psi\!=\!\lambda\psi\quad\Longrightarrow\quad(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\hat{\mathbf{a}}^{\dagger}\psi\!=\!\hat{\mathbf{a}}^{\dagger}(\hat{\mathbf{a}}^{\dagger}\hat{\mathbf{a}}\!+\!\frac{1}{2})\!+\!\hat{\mathbf{a}}^{\dagger}\psi\!=\!(\lambda\!+\!1)\hat{\mathbf{a}}^{\dagger}\psi.$ (78)
Therefore, in the bottom of the spectrum of $\frac{\hat{\mathbf{p}}^{2}\!+\!\hat{\mathbf{q}}^{2}}{2}$, the eigenvalues have an almost equidistant distribution (see Fig. 4).
Figure 4: Eigenvalues of the Hamiltonian (76) in the case d=31d\!=\!31.
Table 7: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied for certain states $\psi_{O,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
IV.4 Discrete coherent states
The discrete coherent states Vourdas04 ; Cotfas11 are
$\begin{array}[]{l}|n,k\rangle\!=\!\frac{1}{\sqrt{d}}\mathbf{D}(n,k)|\mathbf{g}\rangle,\end{array}$ (79)
where $n,k\!\in\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$ and $\mathbf{D}(n,k)\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\begin{array}[]{l}\mathbf{D}(n,k)\psi(m)=\mathrm{e}^{-\frac{\pi\mathrm{i}}{d}nk}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\,\psi(m\!-\!n),\end{array}$ (80)
@@ IV.4 Discrete coherent states
$\parallel\!\mathbf{C}|n,k\rangle\!\parallel^{2}=\!\!\!\sum\limits_{m=-s}^{s}\left|\frac{1}{d}\right.\!\!\sum\limits_{j=-s}^{s}\sum\limits_{\ell=-s}^{s}\!\!j(m\!-\!\ell)\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}j(m-\ell)}\ \mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k\ell}\,\mathbf{g}(\ell\!-\!n)\left.-\,\mathrm{i}\frac{d}{2\pi}\ \mathrm{e}^{\frac{2\pi\mathrm{i}}{d}km}\ \mathbf{g}(m\!-\!n)\right|^{2},$ (81)
@@
$\parallel\!\mathbf{C}|n,k\rangle\!\!\parallel=\parallel\!\mathbf{C}|-n,k\rangle\!\!\parallel=\parallel\!\mathbf{C}|n,-k\rangle\!\!\parallel=\parallel\!\mathbf{C}|-n,-k\rangle\!\!\parallel.$ (82)
@@
Table 8: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation for certain states $|j,k\rangle$.
$|n,k\rangle$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$   $|\ |\langle n,k|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$ $|n,k\rangle$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$ $|\ |\langle n,k|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$
@@
IV.5 Eigenfunctions of discrete annihilation operator
It is known that, in the continuous-variable case, the standard coherent states $|q,p\rangle$ can be defined as eigenstates of the annihilation operator, $\hat{a}|\alpha\rangle\!=\!\alpha|\alpha\rangle$, that is
By following the analogy, we consider the eigenstates $\psi_{A,n}$ of the discrete version
[…]
of $\hat{a}$. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{A,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 9), in the case d=31d\!=\!31, the eigenstates $\psi_{A,0}$, $\psi_{A,1}$, … , $\psi_{A,14}$, where the states have been considered in the increasing order of the number of sign oscillations, belong to $\mathcal{S}_{0.001}$.
h $\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$      $\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$ $\psi_{A,0}$ $2.08125\!\times\!10^{-9}$ 0       $\psi_{A,10}$ $3.2588\!\times\!10^{-6}$ $2.75335\!\times\!10^{-14}$ $\psi_{A,1}$ $2.08125\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$     $\psi_{A,11}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,2}$ $2.08126\!\times\!10^{-9}$ $1.77636\!\times\!10^{-15}$     $\psi_{A,12}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,3}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$     $\psi_{A,13}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,4}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$     $\psi_{A,14}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,5}$ $1.41314\!\times\!10^{-7}$ $1.77636\!\times\!10^{-15}$     $\psi_{A,15}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,6}$ $1.41314\!\times\!10^{-7}$ 0       $\psi_{A,16}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,7}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$     $\psi_{A,17}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,8}$ $3.2588\!\times\!10^{-6}$ $2.66454\!\times\!10^{-14}$     $\psi_{A,18}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,9}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$     $\psi_{A,19}$ 0.0381896 $2.72907\!\times\!10^{-6}$
Table 9: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation by certain states $\psi_{A,k}$.
IV.6 Discrete coherent state quantization
@@
$\begin{array}[]{l}g(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\,\mathrm{e}^{-\frac{1}{2}q^{2}}\end{array}$ (86)
is the vacuum state, satisfy the decomposition of the identity
@@
The map $f\!\mapsto A_{f}$ is called coherent state quantization Cotfas11 ; Gazeau ; Perelomov . One can prove that Gazeau
@@
by following the analogy with the continuous case, one can associate the operator $\mathbf{A}_{f}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\begin{array}[]{l}\mathbf{A}_{f}\!=\!\sum\limits_{n,k=-s}^{s}f(n,k)\,|n,k\rangle\langle n,k|\end{array}$ (91)
@@
defined on the discrete phase space. By using the discrete coherent state quantization $f\!\mapsto\!\mathbf{A}_{f}$, we obtain for $f(n,k)\!=\!\frac{1}{2}(n^{2}\!+\!k^{2})$ the operator
$\begin{array}[]{l}\mathbf{H}_{q}\!=\!-\frac{1}{2}\!+\!\sum\limits_{n,k=-s}^{s}\frac{n^{2}\!+\!k^{2}}{2}\,|n,k\rangle\langle n,k|\end{array}$ (93)
which can be regarded as a discrete version of the harmonic oscillator Hamiltonian. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{Q,n}$ of $\mathbf{H}_{q}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 10), in the case d=31d\!=\!31, the states $\psi_{Q,0}$, $\psi_{Q,1}$, … , $\psi_{Q,7}$, where the states have been considered in the increasing order of eigenvalues, belong to $\mathcal{S}_{0.001}$. One can again remark an almost equidistant distribution in the bottom of the spectrum (see Fig. 5). The Hamiltonian (93) is an other discrete version of (77), and a behavior similar to (76) is again expected to happen.
Table 10: In the case d=31d\!=\!31, the relations $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied with approximation of certain states $\psi_{Q,k}$.
$\psi$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$   $|\ |\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
@@
Figure 5: Eigenvalues of the Hamiltonian (93) in the case d=31d\!=\!31.
IV.7 Mehta states